The formula C = (F - 32) expresses the relationship between Fahrenheit temperature, F, and Celsius temperature, C. Use the formula to convert 41°F to its equivalent temperature on the Celsius scale.
1
Expert's answer
2020-09-06T18:06:40-0400

"F=41, C= \frac{5}{9}(F-32)=\frac{5}{9}(41-32)=\frac{5}{9}*9=5
\\41°F=5°C"